An alpha particle has 3.81 MeV of energy. If an elastic collision with an electron occurs, how much energy will the alpha particle lose?

Give the answer in MeV and to 4 decimal places.
Answer #1

KE = 3.81 Mev =6.104 x 10^ -13 J

mass of alpha particle: 6.64424 x 10-27 kg

1/2 mv^ 2= 6.104 x 10^ -13 J ​

0.5 ( 6.64424. 10-27 ) v^2 = 6.104 x 10^ -13 J ​​

initial velocity = 1.355 x 10^7 m/s

mass of electron = 9.109 x 10^-31 kg

mass of alpha/ mass of electron= (6.64424 x 10-27 kg) / ( 9.109 x 10^-31 kg ​) = 7290

v ( velocity of alpha after collision) = ( 6.64424 x 10-27 - 9.109 x 10^-31 kg​ )1.355 x 10^7​ / (  6.64424 x 10-27 + 9.109 x 10^-31 kg​)

v = (6.6433x 10^-27 ) / (6.6451 x 10^-27) ( 1.355 x 10^7) = 1.3456 x 10^7 m/s

Energy of rebound alpha particle = 0.5 (6.64424 x 10-27 kg) ( 1.3456 x 10^7​)^2 = 6.01516 x 10^ -13 J = 3.75436773 Mev

lost = 3.81 -  3.75436773 =0.05563 MeV apprx